The NOx concentration is given as 1.42 x 10-7 g L-1. Convert this concentration to ppm. (hint: assume an average molar mass of the NOx compounds).

Does this concentration meet the NAAQS for the 1-hour average?
Answer #1

Following is the - complete Answer -&- Explanation: for the given Question:L in...Typed format...

Answer:

  1. Concentration of NOx , in ppm = 0.000142 ppm
  2.   YES, it will stand the NAAQS standards for NOx. It will be less than the NAAQS standards set for NOx....  

Explanation:

  • Part - 1:

We know since:  1 ppm = 1 mg/L  , we get: 1 g/L = 1000 mg/L = 1000 ppm

Therefore: 1.42 x 10-7 g/L = ( 1.42 x 10-7 g/L ) x 1000 = 1.42 x 10-4 ppm

  • Answer: Part - 1: Concentration of NOx , in ppm = 0.000142 ppm

​​​​​​​

  • Part - 2:

​​​​​​​Answer: YES, it will stand the NAAQS standards for NOx,

National Ambient Air Quality Standards ( NAAQS ), has set the maximum, 1-hr standard at a level of 100 ppb .

We know, 1.0 ppb = 0.001 ppm

Therefore: 100 ppb = 0.1 ppm > 0.000142 ppm ( i.e. as calculated above ... )

Therefore: we can surely say that the given concentration of NOx  , will stand, the NAAQS standards....